OpenStudy (anonymous):

Write an equation for the line that passes through the points (5, 6) and (-7, -8). A. y = 6/7 x + 12/7 B. y = - x + 11 C. y = x + 1 D. y = 7/6 x + 1/6

OpenStudy (anonymous):

\[y-y1=\frac{ y2-y1 }{ x2-x1 }\left( x-x1 \right)\]

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

can you show me your work ?

OpenStudy (anonymous):

i just guess

OpenStudy (anonymous):

\[y-6=\frac{ -8-6 }{ -7-5 }\left( x-5 \right)\] \[y-6=\frac{ -14 }{ -12 }\left( x-5 \right)=\frac{ 7 }{ 6 }\left( x-6 \right)\] \[y-6=\frac{ 7 }{6 }x-\frac{ 35 }{ 6 },y=\frac{ 7 }{ 6 }x-\frac{ 35 }{6 }+6\] you can simplify it,it is not a.